2. Andrews Company currently has the following balances in their liability and equity accounts:

Total Liabilities: $18,382,509
Common Stock: $13,156,113
Retained Earnings: $49,722,442

Suppose next year the Andrews Company generates $36,500,000 in Net Profit, pays $15,000,000 in dividends, and total liabilities and common stock remain unchanged. What must their total assets be next year?

a. $132,761,064

b. $102,761,064

c. $49,722,442

d. $81,261,064
Answer #1

Addition to retained earnings=Net profit-Dividends

=(36,500,000-15,000,000)=21,500,000

Hence retained earnings at end=retained earnings at beginning+Addition to retained earnings

=(49,722,442+21,500,000)=$71,222,442

Equity would be=Common stock+retained earnings

=(13,156,113+71,222,442)=$84,378,555

Total assets would be=Debt+Equity

=$84,378,555+18,382,509

=$102,761,064